OpenStudy (anonymous):

∫xcsc^2(2x-1) dx please integrate . thank you very much :)

OpenStudy (amistre64):

\[\int f'(u)f(u)du\]

OpenStudy (amistre64):

prolly shoulda used more of a chainrule looking notation shouldnta i

OpenStudy (amistre64):

and its a first thought anyways, usually off by a mile

sam (.sam.):

Integration by parts

OpenStudy (anonymous):

helps if you know that the anti derivative of cosecant squared is -cotangent

OpenStudy (anonymous):

its all about integration of trigonometric functions

OpenStudy (amistre64):

first thought was good lol

OpenStudy (callisto):

∫xcsc^2(2x-1) dx =(-1/2 )∫x d[cot(2x-1)] =(-1/2 ) [x cot(2x-1) - ∫cot(2x-1)dx = (-1/2 ) [x cot(2x-1)] - (-1/2 ) ∫cot(2x-1)dx = (-1/2 ) [x cot(2x-1)] +(1/2 ) ∫cos(2x-1) / sin(2x+1)dx = (-1/2 ) [x cot(2x-1)] + (1/2)(1/2)∫ 1/sin(2x+1) d (sin(2x+1) = (-1/2 ) [x cot(2x-1)] + (1/4) ln|sin(2x+1)| +C

OpenStudy (amistre64):

thats doesnt look quite right

OpenStudy (amistre64):

x csc^2(2x-1) comes from something similar to -cot(2x-1)

OpenStudy (amistre64):

im reading the x wrong ....

OpenStudy (amistre64):

♫♫ ...waking up is hard to do

OpenStudy (callisto):

Was i incorrect?

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+xcsc%5E2%282x-2%29 callisto is right in that this thing is ugly :/

OpenStudy (amistre64):

you sure it aint 2x^2 - 1 ??

OpenStudy (callisto):

what do you mean?

OpenStudy (amistre64):

I mean, is the problem typed correctly? because what we have as a result seems a bit messy to me ....

OpenStudy (callisto):

not very messy..

OpenStudy (amistre64):

im curious about you changing csc^2 to a d[....]; ive never considered that before

OpenStudy (amistre64):

x d[cot(2x-1)] u = x ; du = dx v = ...... ; dv = d/dx cot(2x-1) im not quite sure how that would make a transition

OpenStudy (amistre64):

v = -1/2 cot(2x-1) ; dv = csc^2(2x-1) dx so it look like a fine manuever

OpenStudy (callisto):

hmm.. just put it inside [i meant the d(x)] by integrating it.. Students here just do it like that.. d/dx [cot(2x-1)] = 2csc^2 (2x-1) therefore ∫csc^2 (2x-1) dx = (1/2)cot(2x-1) +C and ∫xcsc^2 (2x-1) dx = (1/2)∫x d[cot(2x-1)] eh.. i don't really know how to explain..